OpenStudy (anonymous):

prove that [1 - e^(1/x)]/ [1+ e^(1/x)] is odd.

OpenStudy (anonymous):

replace x by -x and see that you get the same thing back. would you like me to work out the gory detail?

OpenStudy (anonymous):

please if you could thatd be great

OpenStudy (anonymous):

\[f(x)=\frac{1-e^{\frac{1}{x}}}{1+e^{\frac{1}{x}}}\] \[f(-x)=\frac{1-e^{-\frac{1}{x}}}{1+e^{-\frac{1}{x}}}\] multiply top and bottom by \[e^{\frac{1}{x}}\] to get \[f(-x)=\frac{e^{\frac{1}{x}}-1}{e^{\frac{1}{x}}+1}=-\frac{1-e^{\frac{1}{x}}}{1+e^{\frac{1}{x}}}\] so \[f(-x)=-f(x)\] and it is odd

OpenStudy (anonymous):

if any step is not clear let me know.

OpenStudy (anonymous):

thnxs a lot

OpenStudy (anonymous):

welcome